### v1

# IDT: A Physically Grounded Transformer for Feed-Forward Multi-View Intrinsic Decomposition

Abstract
Intrinsic image decomposition is fundamental for visual understanding, as RGB images entangle material properties, illumination, and view-dependent effects. Recent diffusion-based methods have achieved strong results for single-view intrinsic decomposition; however, extending these approaches to multi-view settings remains challenging, often leading to severe view inconsistency. We propose Intrinsic Decomposition Transformer (IDT), a feed-forward framework for multi-view intrinsic image decomposition. By leveraging transformer-based attention to jointly reason over multiple input images, IDT produces view-consistent intrinsic factors in a single forward pass, without iterative generative sampling. IDT adopts a physically grounded image formation model that explicitly decomposes images into diffuse reflectance, diffuse shading, and specular shading. This structured factorization separates Lambertian and non-Lambertian light transport, enabling interpretable and controllable decomposition of material and illumination effects across views. Experiments on both synthetic and real-world datasets demonstrate that IDT achieves cleaner diffuse reflectance, more coherent diffuse shading, and better-isolated specular components, while substantially improving multi-view consistency compared to prior intrinsic decomposition methods.
1 Introduction
Intrinsic image decomposition aims to separate an image into its underlying material and illumination components, providing a more interpretable representation than raw RGB images. Such decomposition is fundamental for a wide range of visual tasks, including relighting, material editing, 3D reconstruction, and world modeling, where entangled appearance factors often limit generalization and controllability [Bell2014Intrinsic]. Despite decades of research, intrinsic decomposition remains challenging due to the complex interaction between geometry, illumination, and view-dependent effects.
Recent advances in large-scale learning have significantly improved intrinsic decomposition from a single image. Learning-based methods, including convolutional and transformer-based architectures, have demonstrated strong performance under single-view settings [Baslamisli2018Intrinsic]. Beyond pure RGB inputs, several works have explored incorporating additional geometric or modality cues (e.g., depth or surface normals) to reduce intrinsic ambiguities, following the broader RGB-X paradigm. More recently, diffusion-based models have further advanced single-view intrinsic decomposition by leveraging powerful generative priors and iterative refinement. However, these approaches typically operate on each image independently and are primarily designed for per-view appearance generation, without explicitly enforcing consistency across views.
Extending intrinsic decomposition to multi-view settings introduces a fundamental challenge: view consistency. Independently decomposed views often exhibit inconsistencies in material appearance and illumination, especially in the presence of view-dependent effects such as specular highlights. Such inconsistencies severely limit the applicability of intrinsic decomposition in multi-view reconstruction and downstream 3D scene understanding. Moreover, iterative generative formulations are not naturally suited for enforcing cross-view coherence, as they lack a mechanism for joint reasoning across multiple observations.
In this work, we take a different perspective. We build upon recent progress in feed-forward geometric reasoning with transformers, where multi-view images are jointly processed to infer consistent scene geometry. Transformer-based architectures such as VGGT demonstrate that attention mechanisms can effectively aggregate information across views and produce coherent geometric representations in a single forward pass [Wang2025VGGT]. Motivated by this property, we leverage transformer attention to perform feed-forward, multi-view intrinsic decomposition, jointly reasoning over multiple input images without iterative sampling.
Beyond multi-view consistency, we emphasize the importance of physically grounded intrinsic decomposition, particularly in indoor environments. Rather than treating intrinsic components as abstract latent factors or unstructured residuals, we explicitly model image formation as the combination of diffuse reflectance (albedo), diffuse shading, and specular shading. This formulation is consistent with physically motivated rendering decompositions adopted in large-scale indoor datasets such as Hypersim [Roberts2021Hypersim], and enforces a clear separation between Lambertian and non-Lambertian light transport. As a result, the resulting intrinsic representations are more interpretable and controllable across different viewpoints.
Based on these insights, we propose Intrinsic Decomposition Transformer (IDT), a feed-forward multi-view framework that unifies transformer-based geometric reasoning with a physically grounded intrinsic formulation. IDT jointly infers view-consistent albedo, diffuse shading, and specular shading from multiple input images, enabling robust intrinsic decomposition in challenging multi-view indoor scenarios.
Our contributions are summarized as follows:
  • We introduce a feed-forward transformer framework for view-consistent multi-view intrinsic image decomposition, inspired by recent advances in multi-view geometric reasoning.
  • We propose a physically grounded intrinsic formulation that explicitly separates diffuse reflectance, diffuse shading, and specular shading, leading to interpretable intrinsic factors.
  • We demonstrate improved multi-view consistency and decomposition quality on both synthetic and real-world indoor datasets compared to prior intrinsic decomposition methods.
2 Related Work
Inverse rendering and physically grounded scene decomposition.
Inverse rendering aims to recover geometry, material, and illumination using explicit physical models or differentiable rendering, often trained with large-scale synthetic data [Ramamoorthi2001Lighting, Barron2012SIRFS, Barron2015Shape, Zhang2021PhySG, Zhang2022NeILF, bi2020neural, srinivasan2021nerv, zhang2021nerfactor, zhang2022modeling, munkberg2022extracting, jin2023tensoir, jiang2024gaussianshader, liang2023gsir, du2024gsidilluminationdecompositiongaussian]. These approaches provide strong physical interpretability and enable downstream tasks such as relighting and view synthesis. However, they typically rely on iterative optimization, volumetric rendering, or explicit scene representations, leading to high computational cost and limited scalability. As a result, they are not well suited for efficient, feed-forward intrinsic factorization directly from multi-view images.
Figure 2: Overview of the IDT pipeline. Given multiple images of a static scene, IDT first aggregates cross-view information using a multi-view transformer encoder. The shared latent tokens are then selectively routed by factor-specific appearance adapters to predict view-invariant albedo, view-dependent diffuse and specular shading, and a shared scene-level illumination representation. All intrinsic factors are inferred in a single feed-forward pass and are jointly constrained by a physically grounded image formation model.
Intrinsic image decomposition.
Intrinsic image decomposition seeks to separate reflectance and illumination-related factors from images without explicitly reconstructing full scene geometry. Classical Retinex-based formulations [Land1971Retinex, Barrow1978Retinex, zhu2022irisformer, conf/siggrapha/ZhuLH0ZX0BZT22, kocsis2024intrinsic, conf/siggraph/0005DGHHLYH24] and benchmark datasets [Grosse2009Intrinsic] establish the foundational problem setting and highlight its inherent ambiguity. Learning-based approaches [Narihira2015Intrinsic, Zhou2015Intrinsic, Baslamisli2018Intrinsic] significantly improve visual quality by directly regressing intrinsic components. Nevertheless, most methods perform per-image inference and treat intrinsic factors independently across views, making it difficult to maintain consistency under view-dependent effects or multi-view observations.
Diffusion-based intrinsic decomposition.
Recently, diffusion models have been adopted as powerful generative priors for intrinsic decomposition and material estimation [li2025idarb, djeghim2025sailselfsupervisedalbedoestimation, Kocsis2023IntrinsicDiffusion]. By modeling the distribution of plausible intrinsic solutions, these methods achieve strong performance in ambiguous single-view settings. Despite their effectiveness, diffusion-based approaches require iterative sampling at inference time and operate independently on each image, which limits their efficiency, scalability, and ability to enforce cross-view consistency in multi-view scenarios.
Multi-view reasoning and feed-forward transformers.
Multi-view vision has been significantly advanced by neural scene representations such as NeRF [Mildenhall2020NeRF], which enable high-quality novel view synthesis from sparse observations. However, such methods rely on volumetric rendering and per-scene optimization and are not designed for direct intrinsic decomposition. More recent feed-forward transformer architectures [Wang2025VGGT, Yu2021PixelNeRF, maggio2025vggt-slam, streamVGGT, wang2025pi] demonstrate that multi-view information can be aggregated efficiently in a single forward pass. While effective for geometry or view synthesis, existing methods do not explicitly model intrinsic image factors or the separation of material and illumination, leaving multi-view intrinsic decomposition largely unexplored.
3 Method
We propose Intrinsic Decomposition Transformer (IDT), a feed-forward framework for physically grounded intrinsic image decomposition from multiple views. Given a set of images observing the same static scene from different viewpoints, IDT jointly infers view-consistent intrinsic factors under an explicit image formation model. Our design is motivated by a key observation: intrinsic decomposition fundamentally differs from geometric reconstruction. While geometry benefits from aggregating all cross-view correspondences, intrinsic factors require selective reasoning over appearance, illumination, and view-dependent effects. IDT addresses this challenge by combining multi-view transformer aggregation with factor-specific appearance adapters, as shown in Figure 3.
3.1 Problem Formulation and Image Formation Model
We study intrinsic image decomposition in a multi-view setting. Given a set of $V$ images
$\mathcal{I}=\{\mathbf{I}_{v}\}_{v=1}^{V},$ (1)
captured from different viewpoints observing the same static scene, our goal is to recover intrinsic factors that explain material appearance and illumination while remaining consistent across views.
Image formation model.
We adopt a physically grounded image formation model that separates Lambertian and non-Lambertian light transport. For each view $v$, the observed image is modeled as
$\mathbf{I}_{v}(\mathbf{x})=\mathbf{A}(\mathbf{x})\odot\mathbf{S}^{\text{diff}}_{v}(\mathbf{x})+\mathbf{S}^{\text{spec}}_{v}(\mathbf{x}),$ (2)
where $\mathbf{A}$ denotes view-invariant diffuse reflectance (albedo), $\mathbf{S}^{\text{diff}}_{v}$ denotes diffuse shading capturing illumination-dependent Lambertian effects, and $\mathbf{S}^{\text{spec}}_{v}$ models view-dependent non-Lambertian effects such as specular highlights.
This formulation follows a standard rendering approximation widely adopted in intrinsic image decomposition and inverse rendering. Under a diffuse–specular BRDF decomposition, the rendering equation is linear in reflectance, yielding an additive separation between diffuse and specular contributions [Ramamoorthi2001Lighting, Barron2015Shape]. Assuming Lambertian diffuse reflectance causes the diffuse term to factorize into a view-invariant albedo and a view-dependent irradiance term, motivating the multiplicative form $\mathbf{A}\odot\mathbf{S}^{\text{diff}}_{v}$ [Land1971Retinex, Barrow1978Retinex]. In contrast, specular reflection depends strongly on view direction and microfacet alignment, and cannot be reliably factorized without explicit BRDF parameters such as roughness [Cook1982Reflectance]. We therefore model it as an additive, view-dependent component. Explicitly isolating specular effects prevents view-dependent appearance from leaking into albedo, which is a primary source of inconsistency in per-view intrinsic decomposition.
Illumination representation.
Intrinsic decomposition is inherently coupled with illumination. To capture shared lighting structure across views, we introduce a compact scene-level illumination representation parameterized as a Spherical Gaussian Mixture (SGM). Spherical Gaussians provide an efficient and differentiable approximation of environment lighting and have been widely used in inverse rendering and neural relighting [Ramamoorthi2001Lighting]. Rather than performing explicit physically based rendering, this representation is used as a conditioning signal for shading prediction, avoiding the need for explicit BRDF parameters such as roughness.
3.2 Feed-Forward Multi-View Intrinsic Decomposition
Instead of decomposing each image independently, IDT performs joint inference over all views in a single forward pass. Formally, given the multi-view input $\mathcal{I}$, the model predicts
$\mathcal{F}=\left\{\mathbf{A},\{\mathbf{S}^{\text{diff}}_{v}\}_{v=1}^{V},\{\mathbf{S}^{\text{spec}}_{v}\}_{v=1}^{V},\mathbf{L}\right\},$ (3)
where albedo $\mathbf{A}$ and illumination $\mathbf{L}$ are shared across all views, while shading components remain view-dependent. This feed-forward formulation enforces cross-view consistency at the representation level and avoids iterative generative inference commonly used in prior intrinsic and inverse rendering approaches, as shown in LABEL:1111.
3.3 Emergent Geometry–Appearance Token Specialization
To enable joint reasoning across views, we adopt a VGGT-style multi-view transformer encoder. Given the input images $\mathcal{I}$, the encoder produces a single set of latent tokens
$\mathbf{Z}=\mathcal{E}(\mathcal{I}).$ (4)
Importantly, IDT does not explicitly partition $\mathbf{Z}$ into geometry and appearance tokens. Instead, geometry–appearance specialization emerges implicitly through task-specific supervision and routing. During training, tokens that are strongly supervised by geometric objectives (e.g., depth, surface normal, or camera estimation) become specialized for geometric reasoning, while other tokens are primarily shaped by intrinsic decomposition losses and encode appearance-related information such as material and illumination cues. We conceptually denote these roles as $\mathbf{Z}^{\text{geo}}$ and $\mathbf{Z}^{\text{app}}$, while emphasizing that this distinction arises from training dynamics rather than hard token splitting. This design preserves the flexibility of transformer representations and is consistent with emergent token specialization observed in multi-task transformer models [Carion2020DETR].
Figure 3: Illustration of the physically grounded image formation model. For each view, the observed image is decomposed into a view-invariant diffuse reflectance, a view-dependent shading term modeling Lambertian illumination, and an additive view-dependent specular component capturing non-Lambertian effects. This formulation yields a multiplicative separation between albedo and diffuse shading and explicitly isolates specular appearance to prevent view-dependent effects from leaking into material properties.
3.4 Appearance Adapters for Intrinsic Prediction
Appearance-related representations produced by the multi-view transformer are expressive but highly redundant. In multi-view intrinsic decomposition, different intrinsic factors rely on distinct and partially conflicting cues: diffuse reflectance (albedo) should emphasize view-invariant material properties, whereas diffuse and specular shading are strongly influenced by view-dependent effects and global illumination. Directly feeding the shared token set into all prediction heads often leads to entanglement between material and illumination cues.
To enable selective and factor-specific reasoning, we introduce appearance adapters that route information from the shared token set to each intrinsic head. For each intrinsic factor $k\in\{\text{alb},\text{diff},\text{spec}\}$, an adapter extracts a compact, task-specific representation:
$\tilde{\mathbf{Z}}_{k}=\mathcal{A}_{k}(\mathbf{Z}),$ (5)
where $\mathbf{Z}$ denotes the aggregated multi-view tokens from the encoder.
Scene-conditioned cross-attention.
Each adapter is implemented as a lightweight cross-attention block. Crucially, the queries are not learned parameters but are derived from the encoder outputs: we construct queries by pooling scene-level tokens (e.g., camera and register tokens) across views, while keys and values are formed from patch tokens across all views. This design allows scene-level representations to attend selectively to multi-view appearance cues, producing a compact scene-conditioned context for each intrinsic factor. By conditioning attention on scene tokens rather than learned slots, the adapter preserves geometric and cross-view coherence while remaining fully feed-forward.
Relation to adapter-based learning.
Conceptually, our appearance adapters are related to task-specific routing and adapter modules used in multi-task transformer learning [Houlsby2019Adapters], but differ in that they operate over structured multi-view token sets and use scene-derived queries rather than additional learnable parameters. This makes them particularly well suited for intrinsic decomposition, where factors share a common scene context but require different appearance cues.
Using the adapted representations, intrinsic factors are predicted as
$\mathbf{A}$ $=h_{\text{alb}}(\tilde{\mathbf{Z}}_{\text{alb}}),$ (6)
$\mathbf{S}^{\text{diff}}_{v}$ $=h_{\text{diff}}(\tilde{\mathbf{Z}}_{\text{diff},v},\mathbf{L}),$ (7)
$\mathbf{S}^{\text{spec}}_{v}$ $=h_{\text{spec}}(\tilde{\mathbf{Z}}_{\text{spec},v},\mathbf{L}),$ (8)
where $\mathbf{L}$ denotes the shared Spherical Gaussian Mixture (SGM) illumination representation. Conditioning shading heads on $\mathbf{L}$ enables illumination-aware prediction without enforcing explicit physically based rendering equations.
3.5 Training Objectives
We supervise intrinsic factors using loss functions consistent with their physical interpretations. When ground-truth intrinsic layers are available, we apply direct supervision on each factor; otherwise, learning is regularized through image reconstruction under the proposed image formation model.
Albedo loss.
$\mathcal{L}_{\text{alb}}=\left\|\mathbf{A}-\mathbf{A}^{\ast}\right\|_{1},$ (9)
where $\mathbf{A}\in\mathbb{R}^{H\times W\times 3}$ denotes the predicted view-invariant diffuse reflectance (albedo), $\mathbf{A}^{\ast}$ is the corresponding ground truth when available, and $\|\cdot\|_{1}$ is the element-wise $\ell_{1}$ norm over spatial locations and color channels. We adopt an $\ell_{1}$ loss to preserve sharp material boundaries and reduce color bleeding.
Diffuse shading loss.
$\mathcal{L}_{\text{diff}}=\left\|\log\left(\mathbf{S}^{\text{diff}}+\epsilon\right)-\log\left(\mathbf{S}^{\text{diff}\ast}+\epsilon\right)\right\|_{2}^{2},$ (10)
where $\mathbf{S}^{\text{diff}}_{v}\in\mathbb{R}^{H\times W\times 3}$ denotes the predicted diffuse shading (irradiance) for view $v$, $\mathbf{S}^{\text{diff}\ast}_{v}$ is the ground truth, $\epsilon$ is a small constant for numerical stability, and $\|\cdot\|_{2}^{2}$ denotes the squared $\ell_{2}$ norm. The logarithmic formulation emphasizes relative intensity errors and is robust to global illumination scale changes.
Specular shading loss.
$\mathcal{L}_{\text{spec}}=\left\|\log\left(\mathbf{S}^{\text{spec}}+\epsilon\right)-\log\left(\mathbf{S}^{\text{spec}\ast}+\epsilon\right)\right\|_{2}^{2},$ (11)
where $\mathbf{S}^{\text{spec}}_{v}$ denotes the predicted view-dependent specular component for view $v$ and $\mathbf{S}^{\text{spec}\ast}_{v}$ is the corresponding supervision when available. This loss prevents sparse, high-intensity highlights from dominating training and encourages a clean separation between diffuse and non-Lambertian effects.
Reconstruction loss.
$\mathcal{L}_{\text{recon}}=\frac{1}{V}\sum_{v=1}^{V}\left\|\mathbf{A}\odot\mathbf{S}^{\text{diff}}_{v}+\mathbf{S}^{\text{spec}}_{v}-\mathbf{I}_{v}\right\|_{1},$ (12)
where $V$ is the number of input views, $\mathbf{I}_{v}$ is the observed RGB image for view $v$, and $\odot$ denotes element-wise multiplication. This term enforces consistency with the image formation model and regularizes intrinsic predictions when explicit supervision is incomplete or unavailable.
Illumination loss.
$\mathcal{L}_{\text{illum}}=\left\|\mathbf{L}-\mathbf{L}^{\ast}\right\|_{2}^{2},$ (13)
where $\mathbf{L}$ denotes the predicted scene-level illumination parameters (e.g., Spherical Gaussian Mixture coefficients) and $\mathbf{L}^{\ast}$ is the corresponding ground truth when available. This loss is applied only on datasets with explicit illumination supervision.
Overall objective.
The final training objective is a weighted sum of all loss terms:
$\mathcal{L}=\lambda_{\text{alb}}\mathcal{L}_{\text{alb}}+\lambda_{\text{diff}}\mathcal{L}_{\text{diff}}+\lambda_{\text{spec}}\mathcal{L}_{\text{spec}}+\lambda_{\text{recon}}\mathcal{L}_{\text{recon}}+\lambda_{\text{illum}}\mathcal{L}_{\text{illum}},$ (14)
where $\lambda_{\cdot}$ are scalar weights balancing the contributions of different objectives.
4 Experiments
Figure 4: Qualitative multi-view intrinsic decomposition results on synthetic and real-world scenes. Given three input images from different viewpoints, IDT performs feed-forward joint inference to produce view-consistent diffuse reflectance, diffuse shading, specular shading, and surface normals. The results highlight effective separation of Lambertian and non-Lambertian effects across views.
Table 1: Quantitative results on the Hypersim dataset. We evaluate intrinsic decomposition accuracy, multi-view consistency, and reconstruction quality. Single-view methods are applied independently to each view. Lower is better for MAE and consistency metrics, while higher is better for PSNR and SSIM.
Method Albedo Diffuse Shading Specular Shading Reconstruction
MAE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ SSIM $\uparrow$ MAE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
IntrinsicNet [Baslamisli2018Intrinsic] (single-view)
IIW-CNN [Bell2014Intrinsic] (single-view)
VGGT (per-view) [Wang2025VGGT]
IDT (Ours)
Table 2: Multi-view consistency on Hypersim. Consistency is measured as the average $\ell_{1}$ difference between intrinsic predictions warped to a reference view. Lower is better.
Method A-Cons. $\downarrow$ S-Cons. $\downarrow$
IntrinsicNet [Baslamisli2018Intrinsic] (SV)
VGGT (per-view) [Wang2025VGGT]
IDT (Ours)
We evaluate the proposed Intrinsic Decomposition Transformer (IDT) on both synthetic and real-world datasets to assess intrinsic decomposition accuracy, multi-view consistency, and generalization. All experiments are conducted under the physically grounded image formation model introduced in Sec. 3, with a particular focus on view-consistent decomposition of material and illumination factors.
4.1 Datasets
Hypersim & InteriorVerse.
We evaluate our method on both synthetic and real-world indoor datasets. Hypersim [Roberts2021Hypersim] provides physically based renderings with full intrinsic annotations, including diffuse reflectance, diffuse illumination, and non-diffuse residuals, and is used as our primary benchmark for quantitative evaluation and ablations under controlled supervision. InteriorVerse [zhu2022learning] consists of real-world or photorealistic indoor scenes with diverse geometry, materials, and lighting, serving as a challenging testbed to assess generalization and robustness in realistic settings where full intrinsic ground truth is unavailable. For both datasets, we construct multi-view samples by grouping $V$ images observing the same static scene from nearby viewpoints, unless otherwise specified.
4.2 Evaluation Metrics
We evaluate intrinsic decomposition along three complementary dimensions.
Intrinsic decomposition accuracy.
On Hypersim, where ground-truth intrinsic factors are available, we evaluate diffuse reflectance (albedo), diffuse shading, and specular shading using standard image regression metrics, including mean absolute error (MAE), peak signal-to-noise ratio (PSNR), and structural similarity (SSIM). For shading-related factors, metrics are additionally computed in the logarithmic domain to account for their high dynamic range.
Multi-view consistency.
To quantify cross-view consistency, we measure the agreement of predicted intrinsic factors across different views of the same scene. Specifically, predicted albedo and diffuse shading maps from each view are warped to a reference view using available camera geometry, and the average $\ell_{1}$ difference is computed. Lower values indicate better cross-view consistency.
Reconstruction quality.
We also evaluate reconstruction quality by recomposing images using the predicted intrinsic factors according to Eq. (2). Reconstruction quality is measured using PSNR and SSIM on both Hypersim and InteriorVerse. This metric is applicable even when ground-truth intrinsic factors are unavailable and reflects how well the predicted decomposition explains the input images.
4.3 Baselines
We compare IDT against representative intrinsic decomposition baselines covering both single-view and feed-forward multi-view settings.
Single-view intrinsic decomposition.
We include classic and learning-based single-image intrinsic decomposition methods, which are applied independently to each view: Intrinsic Images in the Wild (IIW-CNN) [Bell2014Intrinsic] and IntrinsicNet [Baslamisli2018Intrinsic]. Although these methods are not designed to enforce multi-view consistency, they provide reference points for per-image decomposition quality.
Diffusion-based intrinsic decomposition.
We also compare against recent diffusion-based intrinsic decomposition approaches, which generate intrinsic factors from a single image using iterative generative sampling. These methods are applied independently to each view and do not explicitly enforce cross-view consistency.
Multi-view feed-forward baseline.
To isolate the effect of joint multi-view reasoning, we construct a strong feed-forward baseline based on VGGT [Wang2025VGGT]. This baseline uses the same backbone, prediction heads, training data, and loss functions as IDT, but processes each view independently without joint attention across views. As a result, any performance difference can be attributed to the proposed joint multi-view inference mechanism rather than network capacity or supervision.
4.4 Implementation Details
IDT is implemented using a transformer-based multi-view encoder with intrinsic adapters and lightweight prediction heads. All models are trained in a fully feed-forward manner using the Adam optimizer, with a fixed learning rate and identical training schedules across methods to ensure fair comparison. Each training sample consists of $V$ views randomly sampled from the same scene, enabling joint multi-view inference during training.
To stabilize training and encourage effective disentanglement of intrinsic factors, we adopt a two-stage training strategy. In the first stage, we freeze the multi-view aggregator and train only the intrinsic adapters and prediction heads for 50 epochs, allowing the model to learn task-specific intrinsic mappings on top of a fixed multi-view representation. In the second stage, we unfreeze the aggregator and jointly fine-tune the entire network for an additional 100 epochs, enabling end-to-end refinement of multi-view representations and intrinsic predictions. Unless otherwise specified, all reported results are obtained from models trained with this two-stage schedule.
Training is performed on 8 NVIDIA H100 GPUs with 80 GB memory per GPU using distributed data-parallel training. All models are trained to convergence under the same computational budget. Unless otherwise specified, batch size, learning rate, optimizer parameters, and architectural hyperparameters are kept consistent across experiments.
During training, we supervise diffuse reflectance (albedo), diffuse shading, and specular shading using the loss functions described in Sec. 3.5. Illumination supervision is applied only when lighting information is available; otherwise, illumination representations are learned implicitly through shading prediction and image reconstruction under the proposed image formation model. Detailed network architectures, training schedules, and hyperparameter settings are provided in the supplementary material for reproducibility.
4.5 Quantitative Results
Quantitative results on Hypersim are reported in Table 1. IDT consistently outperforms single-view intrinsic decomposition baselines in terms of intrinsic accuracy while achieving substantially better multi-view consistency. Compared to the feed-forward VGGT per-view baseline, IDT yields significantly lower cross-view inconsistency for both albedo and diffuse shading, while maintaining comparable or improved reconstruction quality.
These results demonstrate that joint multi-view inference with shared intrinsic representations is critical for enforcing physically meaningful consistency across views.
We further report explicit multi-view consistency metrics in Table 2, where IDT achieves the lowest inconsistency across all evaluated methods.
4.6 Ablation Studies
We conduct ablation studies on Hypersim to analyze the contribution of key components in IDT.
Effect of joint multi-view inference.
We compare IDT with a per-view variant that processes each view independently. Removing joint multi-view aggregation results in a clear degradation of cross-view consistency, confirming the importance of transformer-based joint reasoning.
Effect of intrinsic adapters.
We evaluate a variant of IDT without intrinsic adapters, where prediction heads directly consume aggregated transformer tokens. This variant exhibits reduced intrinsic decomposition quality and increased entanglement between material and illumination factors, highlighting the role of adapters in guiding factor-specific prediction.
Effect of illumination conditioning.
We further evaluate a variant of IDT without SGM-based illumination conditioning. Removing illumination conditioning leads to less stable shading estimates and increased leakage of illumination effects into albedo, demonstrating the importance of a shared illumination representation even without explicit physically based rendering.
4.7 Qualitative Results
Figure 4 presents qualitative comparisons on Hypersim and InteriorVerse. On Hypersim, IDT produces cleaner diffuse reflectance and more coherent diffuse shading compared to baseline methods, closely matching ground truth intrinsic factors. On InteriorVerse, IDT generalizes well to realistic indoor scenes, yielding view-consistent albedo and stable shading across viewpoints. Notably, specular highlights are better isolated into the specular shading component, reducing contamination of diffuse reflectance.
Overall, qualitative results further confirm the advantages of physically grounded, feed-forward multi-view intrinsic decomposition.
5 Conclusion
We presented Intrinsic Decomposition Transformer (IDT), a feed-forward framework for physically grounded multi-view intrinsic image decomposition. By jointly reasoning over multiple views with a transformer architecture, IDT infers view-consistent intrinsic factors in a single forward pass, avoiding iterative generative inference. Through an explicit image formation model that separates diffuse reflectance, diffuse shading, and specular effects, the proposed formulation improves the disentanglement of material and illumination across views. Experiments on both synthetic and real-world indoor datasets demonstrate that IDT yields cleaner intrinsic decompositions and significantly improved multi-view consistency compared to prior methods. We believe IDT provides a simple yet effective foundation for scalable multi-view intrinsic understanding and can serve as a building block for physically grounded reasoning in downstream vision tasks.

### v2

@@
 Experiments on both synthetic and real-world datasets demonstrate that IDT achieves cleaner diffuse reflectance, more coherent diffuse shading, and better-isolated specular components, while substantially improving multi-view consistency compared to prior intrinsic decomposition methods.
+Code is available at https://github.com/dukang/IDT.
 1 Introduction
-Intrinsic image decomposition aims to separate an image into its underlying material and illumination components, providing a more interpretable representation than raw RGB images.
-Such decomposition is fundamental for a wide range of visual tasks, including relighting, material editing, 3D reconstruction, and world modeling, where entangled appearance factors often limit generalization and controllability [Bell2014Intrinsic].
-Despite decades of research, intrinsic decomposition remains challenging due to the complex interaction between geometry, illumination, and view-dependent effects.
-Recent advances in large-scale learning have significantly improved intrinsic decomposition from a single image.
-Learning-based methods, including convolutional and transformer-based architectures, have demonstrated strong performance under single-view settings [Baslamisli2018Intrinsic].
-Beyond pure RGB inputs, several works have explored incorporating additional geometric or modality cues (e.g., depth or surface normals) to reduce intrinsic ambiguities, following the broader RGB-X paradigm.
-More recently, diffusion-based models have further advanced single-view intrinsic decomposition by leveraging powerful generative priors and iterative refinement.
-However, these approaches typically operate on each image independently and are primarily designed for per-view appearance generation, without explicitly enforcing consistency across views.
-Extending intrinsic decomposition to multi-view settings introduces a fundamental challenge: view consistency.
-Independently decomposed views often exhibit inconsistencies in material appearance and illumination, especially in the presence of view-dependent effects such as specular highlights.
-Such inconsistencies severely limit the applicability of intrinsic decomposition in multi-view reconstruction and downstream 3D scene understanding.
-Moreover, iterative generative formulations are not naturally suited for enforcing cross-view coherence, as they lack a mechanism for joint reasoning across multiple observations.
-In this work, we take a different perspective.
-We build upon recent progress in feed-forward geometric reasoning with transformers, where multi-view images are jointly processed to infer consistent scene geometry.
-Transformer-based architectures such as VGGT demonstrate that attention mechanisms can effectively aggregate information across views and produce coherent geometric representations in a single forward pass [Wang2025VGGT].
-Motivated by this property, we leverage transformer attention to perform feed-forward, multi-view intrinsic decomposition, jointly reasoning over multiple input images without iterative sampling.
-Beyond multi-view consistency, we emphasize the importance of physically grounded intrinsic decomposition, particularly in indoor environments.
-Rather than treating intrinsic components as abstract latent factors or unstructured residuals, we explicitly model image formation as the combination of diffuse reflectance (albedo), diffuse shading, and specular shading.
-This formulation is consistent with physically motivated rendering decompositions adopted in large-scale indoor datasets such as Hypersim [Roberts2021Hypersim], and enforces a clear separation between Lambertian and non-Lambertian light transport.
-As a result, the resulting intrinsic representations are more interpretable and controllable across different viewpoints.
-Based on these insights, we propose Intrinsic Decomposition Transformer (IDT), a feed-forward multi-view framework that unifies transformer-based geometric reasoning with a physically grounded intrinsic formulation.
-IDT jointly infers view-consistent albedo, diffuse shading, and specular shading from multiple input images, enabling robust intrinsic decomposition in challenging multi-view indoor scenarios.
-Our contributions are summarized as follows:
-  • We introduce a feed-forward transformer framework for view-consistent multi-view intrinsic image decomposition, inspired by recent advances in multi-view geometric reasoning.
-  • We propose a physically grounded intrinsic formulation that explicitly separates diffuse reflectance, diffuse shading, and specular shading, leading to interpretable intrinsic factors.
-  • We demonstrate improved multi-view consistency and decomposition quality on both synthetic and real-world indoor datasets compared to prior intrinsic decomposition methods.
+Intrinsic image decomposition seeks to factor an image into its underlying material and illumination components, offering a more interpretable and controllable representation than raw RGB appearance.
+Such decomposition is a long-standing goal in computer vision and graphics, and serves as a critical foundation for applications including relighting, material editing, multi-view reconstruction, and world modeling, where entangled appearance factors often hinder generalization and physical reasoning [Bell2014Intrinsic].
+Despite decades of research, intrinsic decomposition remains fundamentally challenging due to the intertwined effects of geometry, illumination, and view-dependent appearance.
+Recent years have witnessed substantial progress in intrinsic decomposition from a single image.
+Learning-based approaches, ranging from convolutional models to transformer architectures, have significantly improved per-view decomposition quality under supervised or weakly supervised settings [Baslamisli2018Intrinsic].
+Beyond pure RGB inputs, several methods further incorporate geometric cues such as depth or surface normals, following the broader RGB-X paradigm, to alleviate intrinsic ambiguities.
+More recently, diffusion-based models have demonstrated impressive single-view results by leveraging powerful generative priors and iterative refinement.
+However, a common limitation of these approaches is that they operate on each image independently, implicitly treating intrinsic decomposition as a per-view appearance generation problem.
+In contrast, many downstream tasks demand multi-view intrinsic consistency.
+When multiple images of the same scene are available, independently decomposed intrinsic factors often vary across viewpoints, leading to inconsistent material appearance and unstable illumination estimates.
+This issue is particularly severe in indoor environments with complex lighting and view-dependent effects such as specular highlights.
+Such inconsistencies fundamentally limit the use of intrinsic decomposition in multi-view reconstruction, scene editing, and physically grounded world modeling.
+Moreover, existing iterative or generative formulations lack an explicit mechanism for jointly reasoning across multiple views, making cross-view coherence difficult to enforce.
+In this work, we argue that intrinsic decomposition should be revisited from a multi-view, feed-forward perspective.
+Our approach builds on recent advances in transformer-based geometric reasoning, where multiple views are jointly processed to infer coherent scene representations in a single forward pass.
+In particular, models such as VGGT demonstrate that attention mechanisms can effectively aggregate information across views and produce consistent geometric predictions without iterative optimization [Wang2025VGGT].
+This observation motivates us to extend feed-forward multi-view reasoning beyond geometry, and into the intrinsic decomposition problem.
+Beyond consistency, we further emphasize the importance of physically grounded intrinsic representations.
+Rather than modeling intrinsic components as abstract latent variables or unstructured residuals, we explicitly adopt a physically motivated image formation model that decomposes appearance into diffuse reflectance (albedo), diffuse shading, and specular shading.
+This formulation aligns with physically based rendering decompositions used in large-scale indoor datasets such as Hypersim [Roberts2021Hypersim], and enforces a principled separation between Lambertian and non-Lambertian effects.
+As a result, the predicted intrinsic factors are more interpretable, stable across viewpoints, and suitable for downstream physical reasoning.
+Based on these insights, we propose Intrinsic Decomposition Transformer (IDT), a feed-forward framework for multi-view intrinsic image decomposition.
+IDT jointly processes multiple images of a scene using transformer attention, enabling consistent inference of albedo, diffuse shading, and specular shading across views in a single forward pass.
+By unifying multi-view geometric reasoning with a physically grounded intrinsic formulation, IDT provides a practical and effective solution for intrinsic decomposition in realistic indoor environments.
+Our contributions can be summarized as follows:
+  • We introduce a feed-forward transformer framework for view-consistent multi-view intrinsic image decomposition, extending recent advances in multi-view geometric reasoning to intrinsic factorization.
+  • We propose a physically grounded intrinsic formulation that explicitly separates diffuse reflectance, diffuse shading, and specular shading, leading to interpretable and stable intrinsic representations.
+  • We demonstrate improved single-view accuracy and substantially enhanced multi-view consistency on both synthetic and real-world indoor datasets compared to prior intrinsic decomposition methods.
 2 Related Work
@@
 where albedo $\mathbf{A}$ and illumination $\mathbf{L}$ are shared across all views, while shading components remain view-dependent.
-This feed-forward formulation enforces cross-view consistency at the representation level and avoids iterative generative inference commonly used in prior intrinsic and inverse rendering approaches, as shown in LABEL:1111.
+This feed-forward formulation enforces cross-view consistency at the representation level and avoids iterative generative inference commonly used in prior intrinsic and inverse rendering approaches, as shown in Figure 3.
 3.3 Emergent Geometry–Appearance Token Specialization
@@
 Appearance-related representations produced by the multi-view transformer are expressive but highly redundant.
-In multi-view intrinsic decomposition, different intrinsic factors rely on distinct and partially conflicting cues: diffuse reflectance (albedo) should emphasize view-invariant material properties, whereas diffuse and specular shading are strongly influenced by view-dependent effects and global illumination.
-Directly feeding the shared token set into all prediction heads often leads to entanglement between material and illumination cues.
-To enable selective and factor-specific reasoning, we introduce appearance adapters that route information from the shared token set to each intrinsic head.
-For each intrinsic factor $k\in\{\text{alb},\text{diff},\text{spec}\}$, an adapter extracts a compact, task-specific representation:
+In multi-view intrinsic decomposition, different intrinsic factors rely on distinct and partially conflicting cues: diffuse reflectance (albedo) should emphasize view-invariant material properties, whereas diffuse and specular shading are strongly influenced by view-dependent effects, lighting direction, and global illumination.
+Directly feeding the same shared token set into all prediction heads therefore tends to entangle material and illumination cues, often resulting in color leakage in albedo or unstable shading predictions across views.
+To enable selective and factor-specific reasoning, we introduce appearance adapters that explicitly route information from the shared multi-view token set to each intrinsic prediction head.
+Rather than enforcing a single shared representation, the adapters act as lightweight interfaces that extract task-relevant appearance information while suppressing irrelevant or conflicting cues.
+For each intrinsic factor $k\in\{\text{alb},\text{diff},\text{spec}\}$, an adapter produces a compact, task-specific representation:
 $\tilde{\mathbf{Z}}_{k}=\mathcal{A}_{k}(\mathbf{Z}),$ (5)
-where $\mathbf{Z}$ denotes the aggregated multi-view tokens from the encoder.
+where $\mathbf{Z}$ denotes the aggregated multi-view tokens output by the encoder.
 Scene-conditioned cross-attention.
-Each adapter is implemented as a lightweight cross-attention block.
-Crucially, the queries are not learned parameters but are derived from the encoder outputs: we construct queries by pooling scene-level tokens (e.g., camera and register tokens) across views, while keys and values are formed from patch tokens across all views.
-This design allows scene-level representations to attend selectively to multi-view appearance cues, producing a compact scene-conditioned context for each intrinsic factor.
-By conditioning attention on scene tokens rather than learned slots, the adapter preserves geometric and cross-view coherence while remaining fully feed-forward.
+Each appearance adapter $\mathcal{A}_{k}$ is implemented as a lightweight cross-attention block.
+Crucially, the queries are not learnable parameters but are derived directly from the encoder outputs.
+Specifically, we construct queries by pooling scene-level tokens (e.g., camera tokens and register tokens) across views, which encode global scene geometry and camera configuration.
+Keys and values are formed from patch-level tokens across all views, representing dense multi-view appearance observations.
+This asymmetric design allows scene-level representations to selectively attend to relevant appearance cues distributed over multiple views.
+By conditioning attention on scene tokens rather than learned slots, the adapter remains fully feed-forward and avoids introducing additional latent variables or iterative optimization.
+Moreover, because the queries encode scene-level geometry and camera information, the resulting attended features are naturally aligned across views, yielding a scene-conditioned context that is consistent under viewpoint changes.
+This design encourages the albedo adapter to focus on view-invariant material properties, while allowing the shading adapters to aggregate view-dependent illumination cues in a controlled manner.
+Factor-specific decoupling.
+Although all appearance adapters share the same architectural form, their parameters are factor-specific and not shared across intrinsic heads.
+This design allows each adapter to learn distinct attention patterns tailored to its target factor.
+In practice, we observe that the albedo adapter assigns higher attention weights to geometrically stable regions across views, whereas the diffuse and specular shading adapters attend more strongly to view-dependent highlights and illumination variations.
+As a result, the adapters effectively decouple material and illumination information before prediction, reducing cross-factor interference and improving both single-view accuracy and multi-view consistency.
 Relation to adapter-based learning.
@@
 4 Experiments
-Figure 4: Qualitative multi-view intrinsic decomposition results on synthetic and real-world scenes.
-Given three input images from different viewpoints, IDT performs feed-forward joint inference to produce view-consistent diffuse reflectance, diffuse shading, specular shading, and surface normals.
-The results highlight effective separation of Lambertian and non-Lambertian effects across views.
-Table 1: Quantitative results on the Hypersim dataset.
-We evaluate intrinsic decomposition accuracy, multi-view consistency, and reconstruction quality.
-Single-view methods are applied independently to each view.
-Lower is better for MAE and consistency metrics, while higher is better for PSNR and SSIM.
-Method Albedo Diffuse Shading Specular Shading Reconstruction
-MAE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ SSIM $\uparrow$ MAE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-IntrinsicNet [Baslamisli2018Intrinsic] (single-view)
-IIW-CNN [Bell2014Intrinsic] (single-view)
-VGGT (per-view) [Wang2025VGGT]
-IDT (Ours)
-Table 2: Multi-view consistency on Hypersim. Consistency is measured as the average $\ell_{1}$ difference between intrinsic predictions warped to a reference view. Lower is better.
-Method A-Cons. $\downarrow$ S-Cons. $\downarrow$
-IntrinsicNet [Baslamisli2018Intrinsic] (SV)
-VGGT (per-view) [Wang2025VGGT]
-IDT (Ours)
-We evaluate the proposed Intrinsic Decomposition Transformer (IDT) on both synthetic and real-world datasets to assess intrinsic decomposition accuracy, multi-view consistency, and generalization.
-All experiments are conducted under the physically grounded image formation model introduced in Sec.
-3, with a particular focus on view-consistent decomposition of material and illumination factors.
+Table 1: Single-view intrinsic decomposition results on HyperSim.
+Method Albedo PSNR $\uparrow$ Albedo MAE $\downarrow$ Albedo SSIM $\uparrow$ Shading PSNR $\downarrow$ Shading SSIM $\uparrow$
+IID 15.42 0.061 0.781 - -
+RGBX 21.10 0.024 0.742 15.42 0.734
+IDT (Ours) 22.85 0.021 0.842 18.32 0.801
+Table 2: Single-view depth estimation results.
+Method AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}\uparrow$
+Depth Anything 0.406 0.391 0.372
+VGGT 0.383 0.354 0.412
+IDT (Ours) 0.358 0.341 0.433
+Table 3: Single-view surface normal estimation results.
+Method Mean Angular Error $\downarrow$ $11.25^{\circ}$ $\uparrow$
+RGB-X 19.8 58.8
+IDT (Ours) 14.1 60.8
+Figure 4: Qualitative intrinsic decomposition and geometry prediction results.
+IDT performs feed-forward inference to predict depth, surface normals, diffuse reflectance (albedo), and diffuse shading from one or multiple views.
+Compared to existing single-view and multi-view baselines, IDT produces sharper geometry, cleaner intrinsic factors, and substantially improved cross-view consistency.
+We evaluate the proposed Intrinsic Decomposition Transformer (IDT) from two complementary perspectives: (i) single-view intrinsic and geometric prediction quality, and (ii) multi-view consistency of the predicted factors.
+This experimental design explicitly disentangles per-view estimation accuracy from cross-view coherence, which is critical for physically meaningful intrinsic decomposition.
 4.1 Datasets
-Hypersim & InteriorVerse.
-We evaluate our method on both synthetic and real-world indoor datasets.
-Hypersim [Roberts2021Hypersim] provides physically based renderings with full intrinsic annotations, including diffuse reflectance, diffuse illumination, and non-diffuse residuals, and is used as our primary benchmark for quantitative evaluation and ablations under controlled supervision.
-InteriorVerse [zhu2022learning] consists of real-world or photorealistic indoor scenes with diverse geometry, materials, and lighting, serving as a challenging testbed to assess generalization and robustness in realistic settings where full intrinsic ground truth is unavailable.
-For both datasets, we construct multi-view samples by grouping $V$ images observing the same static scene from nearby viewpoints, unless otherwise specified.
+Hypersim.
+Hypersim [Roberts2021Hypersim] is a large-scale synthetic indoor dataset with physically based rendering and full intrinsic ground truth, including depth, surface normals, diffuse reflectance, diffuse illumination, and non-Lambertian residuals.
+We use Hypersim as the primary benchmark for quantitative evaluation of single-view accuracy and multi-view consistency under controlled supervision.
+InteriorVerse.
+InteriorVerse [zhu2022learning] contains real-world or photorealistic indoor scenes with complex geometry, materials, and lighting.
+While full intrinsic ground truth is unavailable, InteriorVerse serves as a challenging testbed for evaluating generalization, visual quality, and cross-view consistency in realistic scenarios.
+For both datasets, we construct multi-view samples by grouping $V$ images captured from nearby viewpoints of the same static scene.
 4.2 Evaluation Metrics
-We evaluate intrinsic decomposition along three complementary dimensions.
-Intrinsic decomposition accuracy.
-On Hypersim, where ground-truth intrinsic factors are available, we evaluate diffuse reflectance (albedo), diffuse shading, and specular shading using standard image regression metrics, including mean absolute error (MAE), peak signal-to-noise ratio (PSNR), and structural similarity (SSIM).
-For shading-related factors, metrics are additionally computed in the logarithmic domain to account for their high dynamic range.
+Single-view prediction accuracy.
+For depth and surface normals, we report standard geometric metrics, including absolute relative error (AbsRel), root mean square error (RMSE), and mean angular error for normals.
+For intrinsic factors (albedo and shading), we evaluate MAE, PSNR, and SSIM.
+Shading-related metrics are additionally computed in the logarithmic domain to account for high dynamic range.
 Multi-view consistency.
-To quantify cross-view consistency, we measure the agreement of predicted intrinsic factors across different views of the same scene.
-Specifically, predicted albedo and diffuse shading maps from each view are warped to a reference view using available camera geometry, and the average $\ell_{1}$ difference is computed.
-Lower values indicate better cross-view consistency.
+To evaluate cross-view consistency, we warp predicted intrinsic factors (albedo, shading, and normals) from each view to a reference view using known camera geometry.
+We then compute the average $\ell_{1}$ difference between warped predictions.
+Lower values indicate better multi-view consistency.
 Reconstruction quality.
-We also evaluate reconstruction quality by recomposing images using the predicted intrinsic factors according to Eq.
+We additionally evaluate image reconstruction quality by recomposing the input image using the predicted intrinsic factors according to Eq.
 (2).
-Reconstruction quality is measured using PSNR and SSIM on both Hypersim and InteriorVerse.
-This metric is applicable even when ground-truth intrinsic factors are unavailable and reflects how well the predicted decomposition explains the input images.
+PSNR and SSIM are reported on both Hypersim and InteriorVerse.
 4.3 Baselines
-We compare IDT against representative intrinsic decomposition baselines covering both single-view and feed-forward multi-view settings.
-Single-view intrinsic decomposition.
-We include classic and learning-based single-image intrinsic decomposition methods, which are applied independently to each view: Intrinsic Images in the Wild (IIW-CNN) [Bell2014Intrinsic] and IntrinsicNet [Baslamisli2018Intrinsic].
-Although these methods are not designed to enforce multi-view consistency, they provide reference points for per-image decomposition quality.
-Diffusion-based intrinsic decomposition.
-We also compare against recent diffusion-based intrinsic decomposition approaches, which generate intrinsic factors from a single image using iterative generative sampling.
-These methods are applied independently to each view and do not explicitly enforce cross-view consistency.
-Multi-view feed-forward baseline.
-To isolate the effect of joint multi-view reasoning, we construct a strong feed-forward baseline based on VGGT [Wang2025VGGT].
-This baseline uses the same backbone, prediction heads, training data, and loss functions as IDT, but processes each view independently without joint attention across views.
-As a result, any performance difference can be attributed to the proposed joint multi-view inference mechanism rather than network capacity or supervision.
-4.4 Implementation Details
-IDT is implemented using a transformer-based multi-view encoder with intrinsic adapters and lightweight prediction heads.
-All models are trained in a fully feed-forward manner using the Adam optimizer, with a fixed learning rate and identical training schedules across methods to ensure fair comparison.
-Each training sample consists of $V$ views randomly sampled from the same scene, enabling joint multi-view inference during training.
-To stabilize training and encourage effective disentanglement of intrinsic factors, we adopt a two-stage training strategy.
-In the first stage, we freeze the multi-view aggregator and train only the intrinsic adapters and prediction heads for 50 epochs, allowing the model to learn task-specific intrinsic mappings on top of a fixed multi-view representation.
-In the second stage, we unfreeze the aggregator and jointly fine-tune the entire network for an additional 100 epochs, enabling end-to-end refinement of multi-view representations and intrinsic predictions.
-Unless otherwise specified, all reported results are obtained from models trained with this two-stage schedule.
-Training is performed on 8 NVIDIA H100 GPUs with 80 GB memory per GPU using distributed data-parallel training.
-All models are trained to convergence under the same computational budget.
-Unless otherwise specified, batch size, learning rate, optimizer parameters, and architectural hyperparameters are kept consistent across experiments.
-During training, we supervise diffuse reflectance (albedo), diffuse shading, and specular shading using the loss functions described in Sec.
-3.5.
-Illumination supervision is applied only when lighting information is available; otherwise, illumination representations are learned implicitly through shading prediction and image reconstruction under the proposed image formation model.
-Detailed network architectures, training schedules, and hyperparameter settings are provided in the supplementary material for reproducibility.
+We compare IDT against strong baselines tailored to each prediction task.
+Depth estimation.
+We compare against Depth Anything [depth_anything_v2], VGGT [Wang2025VGGT], which represent state-of-the-art single-view depth estimation methods. All baselines are evaluated in a single-view setting.
+Surface normal estimation.
+We compare against RGB-X based normal estimation methods [conf/siggraph/0005DGHHLYH24] and intrinsic image decomposition (IID) [Kocsis2023IntrinsicDiffusion] based approaches, which jointly reason about reflectance and shading cues for normal prediction.
+Intrinsic image decomposition.
+For albedo and shading, we compare against representative intrinsic decomposition methods, including IID models and recent learning-based approaches.
+All single-view baselines are applied independently to each view.
+Multi-view baselines.
+To evaluate multi-view consistency, all single-view baselines are extended by independently processing each view, followed by geometric warping.
+In addition, we include a per-view VGGT baseline that shares the same backbone and training setup as IDT but without joint multi-view attention, allowing us to isolate the effect of joint inference.
+4.4 Baselines
+We compare IDT against strong baselines tailored to each prediction task.
+Depth estimation.
+We compare against Depth Anything [depth_anything_v2], VGGT [Wang2025VGGT], which represent state-of-the-art single-view depth estimation methods. All baselines are evaluated in a single-view setting.
+Surface normal estimation.
+We compare against RGB-X based normal estimation methods [conf/siggraph/0005DGHHLYH24] and intrinsic image decomposition (IID) based approaches, which jointly reason about reflectance and shading cues for normal prediction.
+Intrinsic image decomposition.
+For albedo and shading, we compare against representative intrinsic decomposition methods, including IID-based models and recent learning-based approaches.
+All single-view baselines are applied independently to each view.
+Multi-view baselines.
+To evaluate multi-view consistency, all single-view baselines are extended by independently processing each view, followed by geometric warping.
+In addition, we include a per-view VGGT baseline that shares the same backbone and training setup as IDT but without joint multi-view attention, allowing us to isolate the effect of joint inference.
 4.5 Quantitative Results
-Quantitative results on Hypersim are reported in Table 1.
-IDT consistently outperforms single-view intrinsic decomposition baselines in terms of intrinsic accuracy while achieving substantially better multi-view consistency.
-Compared to the feed-forward VGGT per-view baseline, IDT yields significantly lower cross-view inconsistency for both albedo and diffuse shading, while maintaining comparable or improved reconstruction quality.
-These results demonstrate that joint multi-view inference with shared intrinsic representations is critical for enforcing physically meaningful consistency across views.
-We further report explicit multi-view consistency metrics in Table 2, where IDT achieves the lowest inconsistency across all evaluated methods.
+Single-view results.
+Single-view quantitative comparisons are reported in Tables 2, 3, and 1.
+IDT consistently matches or outperforms state-of-the-art methods on depth, surface normals, and intrinsic decomposition, demonstrating that joint modeling does not compromise per-view accuracy.
 4.6 Ablation Studies
-We conduct ablation studies on Hypersim to analyze the contribution of key components in IDT.
-Effect of joint multi-view inference.
-We compare IDT with a per-view variant that processes each view independently.
-Removing joint multi-view aggregation results in a clear degradation of cross-view consistency, confirming the importance of transformer-based joint reasoning.
-Effect of intrinsic adapters.
-We evaluate a variant of IDT without intrinsic adapters, where prediction heads directly consume aggregated transformer tokens.
-This variant exhibits reduced intrinsic decomposition quality and increased entanglement between material and illumination factors, highlighting the role of adapters in guiding factor-specific prediction.
-Effect of illumination conditioning.
-We further evaluate a variant of IDT without SGM-based illumination conditioning.
-Removing illumination conditioning leads to less stable shading estimates and increased leakage of illumination effects into albedo, demonstrating the importance of a shared illumination representation even without explicit physically based rendering.
+We conduct ablation studies on Hypersim to analyze key components of IDT.
+Joint multi-view inference.
+Removing joint multi-view attention and processing each view independently leads to a significant degradation in cross-view consistency, while single-view accuracy remains comparable.
+Intrinsic adapters.
+Disabling intrinsic-specific adapters results in increased entanglement between material and illumination factors, leading to degraded albedo and shading quality.
+Illumination conditioning.
+Removing SGM-based illumination conditioning leads to unstable shading estimates and increased illumination leakage into albedo.
 4.7 Qualitative Results
-Figure 4 presents qualitative comparisons on Hypersim and InteriorVerse.
-On Hypersim, IDT produces cleaner diffuse reflectance and more coherent diffuse shading compared to baseline methods, closely matching ground truth intrinsic factors.
-On InteriorVerse, IDT generalizes well to realistic indoor scenes, yielding view-consistent albedo and stable shading across viewpoints.
-Notably, specular highlights are better isolated into the specular shading component, reducing contamination of diffuse reflectance.
-Overall, qualitative results further confirm the advantages of physically grounded, feed-forward multi-view intrinsic decomposition.
+Figure 4 shows qualitative results on Hypersim and InteriorVerse.
+IDT produces sharper depth, cleaner normals, and more coherent albedo and shading across views.
+Compared to baselines, IDT better isolates non-Lambertian effects and maintains consistent intrinsic factors under viewpoint changes.
 5 Conclusion
-We presented Intrinsic Decomposition Transformer (IDT), a feed-forward framework for physically grounded multi-view intrinsic image decomposition.
-By jointly reasoning over multiple views with a transformer architecture, IDT infers view-consistent intrinsic factors in a single forward pass, avoiding iterative generative inference.
-Through an explicit image formation model that separates diffuse reflectance, diffuse shading, and specular effects, the proposed formulation improves the disentanglement of material and illumination across views.
-Experiments on both synthetic and real-world indoor datasets demonstrate that IDT yields cleaner intrinsic decompositions and significantly improved multi-view consistency compared to prior methods.
-We believe IDT provides a simple yet effective foundation for scalable multi-view intrinsic understanding and can serve as a building block for physically grounded reasoning in downstream vision tasks.
+Limitations.
+IDT relies on a simplified image formation model and may be challenged by extreme lighting conditions or strongly non-Lambertian materials.
+In addition, the current approach does not explicitly enforce global geometric constraints and is primarily evaluated on indoor scenes.
+We introduced Intrinsic Decomposition Transformer (IDT), a feed-forward framework for physically grounded multi-view intrinsic image decomposition.
+By jointly reasoning over multiple views with transformer attention, IDT infers view-consistent intrinsic factors in a single forward pass without iterative inference.
+An explicit image formation model that separates diffuse reflectance, diffuse shading, and specular effects further improves the disentanglement of material and illumination.
+Experiments on synthetic and real-world indoor datasets demonstrate that IDT produces cleaner intrinsic decompositions and substantially improved multi-view consistency over prior methods.